OpenStudy (anonymous):

Need a little help please? At West Painting, they get about three calls a day asking for an estimate of the cost for having the interior of a house painted. To write up an estimate for the cost of a job, they need to know how much paint a job will take. If they average painting 3/4 of a room with 2/5 of a gallon of paint, then they can paint, on average, ____ rooms per gallon.

OpenStudy (anonymous):

@CausticSyndicalist @Conqueror @cj49 @Coolsector @danish071996 @Firefly007 @DanteFemboy @eliassaab @AnswerMyQuestions @sammixboo

OpenStudy (anonymous):

i suck at math so much i am so sorry but i thinks its 4

OpenStudy (anonymous):

ok ur the only one who answer and didnt just sit there so ur appreciated

OpenStudy (anonymous):

i only think that cause, 3/4, 4-3=1, 2/5, 5-2=3, 3+1=4

OpenStudy (anonymous):

thank you, lol when it comes to math im abit daft

OpenStudy (anonymous):

well tht didnt go over my head... ._. anyway thanks

OpenStudy (anonymous):

it was 1 7/8

OpenStudy (anonymous):

:/ sorry